Solve 𝑦 β€²β€² βˆ’ π‘₯𝑦 β€² + 𝑦 = 0 assuming a power series. Find the range of π‘₯ for which the solution is valid


1
Expert's answer
2021-11-15T08:34:34-0500

"y''-xy'+y=0"

It is clear from the equation above that it has no singular point. Therefore the range of x for which the solution is valid is "(-\infty,\infty)"


Let the solution of our equation be of the form "y=\displaystyle\sum_{n=0}^{\infty}a_nx^n" .....(1)

From which,

"y'=\displaystyle\sum_{n=0}^{\infty}na_nx^{n-1}=\displaystyle\sum_{n=1}^{\infty}na_nx^{n-1}"

"y''=\displaystyle\sum_{n=0}^{\infty}n(n-1)a_nx^{n-2}=\displaystyle\sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}"

We then substitute the above expressions in our equation

"\displaystyle\sum_{n=2}^{\infty}n(n-1)a_nx^{n-2}-\displaystyle\sum_{n=1}^{\infty}na_nx^{n}+\displaystyle\sum_{n=0}^{\infty}a_nx^n=0"

"\displaystyle\sum_{n=0}^{\infty}(n+1)(n+2)a_{n+2}x^{n}-\displaystyle\sum_{n=1}^{\infty}na_nx^n+\displaystyle\sum_{n=1}^{\infty}a_nx^n+a_0=0"

"2a_2+6a_3x+\displaystyle\sum_{n=1}^{\infty}[(n+1)(n+2)a_{n+2}-na_n+a_n]x^n+a_0=0"

Comparing both sides of equal powers of x we have,

"2a_2+a_0=0\implies a_2=-\frac{1}{2}a_0"

"6a_3=0\implies a_3=0"

"a_{n+2}=\frac{n-1}{(n+1)(n+2)}a_n" , "n\ge 0" .....(1)

From the recurrence relation (2) above we get,

"a_{4}=\frac{1}{(2+1)(2+2)}a_2=-\frac{1}{24}a_0"

"a_5=0=a_7=a_9=0..."

"a_{6}=\frac{4-1}{(4+1)(4+2)}a_4=\frac{1}{10}\cdot (-\frac{1}{24}a_0)=-\frac{1}{240}a_0"

Substitute the values of "a_2,a_3,a_4,a_5,a_6" in the relation (1) to get:

"y=a_1x+a_0(1-\frac{1}{2}x^2-\frac{1}{24}x^4-\frac{1}{240}x^6...)"

"\therefore" the solution to our equation is

"y=Ax+B(1-\frac{1}{2}x^2-\frac{1}{24}x^4-\frac{1}{240}x^6...)"

Where A and B are arbitrary constants.